

********************************************************************
UARS WEEKLY STATUS REPORT - 02 JUN 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT and
OA_SAT_ATT services. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM AXIS measured X-rays from half of its camera (8 pixels) from
May 3 to May 6 and again from May 20 to May 26. The full AXIS instrument
measured X-rays between May 6 and May 12. At the present time, PEM energy
deposition is not being generated due to failures in the inversion code.
These are currently being examined.
The PEM spectrometers continue to be active under PMON control*. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct for new
telemetry requirements, PEM will need to update processing code with a patch
delivery. Updated HEPS calibration parameters were finalized and are ready for
delivery. Updated MEPS quality calibration files were finalized and are ready
for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell, have
been fixed; his suggestions have been incorporated with in the PEM production
code. Both updated HEPS and MEPS calibration data, along with the Level 3
code changes, will influence the in-situ energy deposition predictions, so new
versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII's last operating sequence was May 19-30, and operations
were normal. The next sequence planned is for June 18-26, when
the objective will be to observe Polar Mesospheric Clouds in
the northern polar summer.
***********
MPG
***********
The MPG attended the UARS Science Team Meeting held this week
in conjunction with the AGU meeting in Washington, DC.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to assist the Project Scientist in
scheduling of instrument operations.
***********
CDHF
***********
CDHF participated in CSOC ISO AUDIT 24-MAY-2000.
Data received and cataloged at the CDHF in the month of March 2000 has been
loaded into the DAAC Transfer Tables for HALOE, NMC, and UKMO.
Confirmation received from DAAC that MLS V4 LEVEL 3AT, ICE_SATURAT (UTH)
UARS DAY 7 - 2103 were all successfully ingested and archived into the
DAAC System 23-May-2000. The data is available for users to order.
The CDHF will load April 2000 current day DAAC data in June 2000.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, MLS,
SUSIM, Solstice, WINDII and HRDI. This completes current day processing
suites that run on ALPHA. HRDI will be notified of location of test results
for validation purposes.
- In the next two weeks we will be doing system test (l0,nrt,production
processing, daac xfers, rac xfers, correlative ingests etc).
***********
FOT
***********
REPORTING PERIOD
This report covers 20 May 2000 (Orbit 47492, GMT Day 141) through 26 May 2000
(Orbit 47597, GMT Day 147).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 141/0011 47492 AXIS-1 On
HRDI 141/1433 47501 On
HALOE 144/0018 47537 On
The beta angle ranged between the angles of 67.3 degrees to 39.7 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+0.0 degrees on 05 June 2000 (DOY 157).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.7 to 1.0 degrees C. Additional battery performance data for each
battery on 19 May 2000 and 26 May 2000 are provided in Appendix D.
The clock error ranged from -10.8 to 11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 16
% HALOE = 01
% HRDI = 01
% PEM = 01
% SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/146 show the observatory data loss to be 295 hours, 38
minutes, 10.561 seconds (an increase of 0 minutes 0.00 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to test the new TMONs on the simulator. Once the TMONs have
been thoroughly tested, they will be presented for review and approval before
uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for June 05, 2000 (DOY 157). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 May 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47492 141/0011 00-224 PEM AXIS 1 Power ON
47496 141/0654 00-217 NBTR PLAYBACK
47501 141/1433 00-223 SUSIM SUSIM CMD
47501 141/1433 00-225 HRDI Power ON FULL-TIME
47511 142/0603 00-217 NBTR PLAYBACK
47521 142/2232 00-226 OBC Clk Adj 20.6 to 0.00 msec/day
47533 143/1759 00-227 ATC NBTR NOOPS D129
47537 144/0018 00-228 HALOE Power ON
47543 144/0850 00-227 NBTR PLAYBACK/NOOPS
47548 144/1801 00-229 NBTR NBTR NOOPS
47548 144/1801 00-230 OBC Clk Adj 0.00 to 20.6 msec/day
47557 145/0712 00-229 NBTR NBTR NOOPS D145
47558 145/0854 00-229 NBTR PLAYBACK
47565 145/2055 00-231 NBTR NBTR NOOPS D146
47571 146/0608 00-231 NBTR NBTR NOOPS D146
47573 146/0857 00-231 NBTR PLAYBACK
47573 146/0937 00-232 OBC Clk Adj 20.6 to 0.00 msec/day
47578 146/1805 00-233 NBTR NBTR NOOPS D147
47586 147/0606 00-233 NBTR NBTR NOOPS D147
47593 147/1807 00-234 NBTR NBTR NOOPS D148
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 May 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTMP Y-HI INTERFEROMETER TEMP 47494-5
HRRADTEMP Y-HI RADIATOR TEMP 47495
HRPTSBARTEMP Y-LO HRDI BARREL TEMP 47492-506
SOSGTEVLTG Y-HI SOLS G-TUBE VOLTAGE 47516
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
SSCMDERR SSPP CMD ERROR 47502
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 May 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47521 142/22:35 (21 May 00) from 20.6 to 0.00 msec/day
47548 144/18:20 (23 May 00) from 0.00 to 20.6 msec/day
47573 146/09:40 (25 May 00) from 20.6 to 0.00 msec/day
______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 May 2000
19 May 2000 - GMT Day 140 Beta = 67.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------ ------- ------
1 +11.2/-5.6 -3.68/-4.62 +0.0/+0.0 15.8 0.000 0.0
2 +16.8/-11.2 +3.87/+2.92 +23.2/-14.4 27.5 1.070 16.5
3 +16.8/-22.4 +2.92/+1.97 +22.4/-13.6 27.5 1.068 16.0
26 May 2000 - GMT Day 147 Beta = 39.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +16.8/+0.0 -3.05/-4.31 +0.0/+0.0 15.7 0.000 0.0
2 +39.2/-33.6 +5.79/+4.51 +28.4/-15.2 24.6 1.030 21.8
3 +22.4/-44.8 +5.15/+3.55 +27.6/-16.0 24.6 1.028 21.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 May 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-021 FOT 47510 142/0514 14170 Generic Late Acq
00-021 FOT 47588 147/0900 14170 Generic Late Acq
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list May 00 14170 Generic Late Acq.
SEVNTFW00/21